A sphere has a volume of 25m^3. What is the diameter in dm??
Answer by stanbon(75887)

A sphere has a volume of 25m^3. What is the diameter in dm??
------------
V = (4/3)pi*r^3
----
25 = (4/3)pi*r^3
75/(4pi) = r^3
r^3 = 5.9683
r = 1.8139
2r = 3.6278
-
diameter = 3.6278 meters
---
diameter = 3.6278(meters)(10 decimeters/1 meter) = 36.278 decimeters
=======================
Cheers,
Stan H.